1. The Andoh Family buys a house for GHȼ94,000 with a down payment of GHȼ16,000. They

take out a 30-year mortgage for GHȼ78,000 at an annual interest rate of 9.6%.

a. Find the amount of the monthly payment to amortize this loan.

b. Find the total amount of interest paid when the loan is amortized after 30years.

2. Compute the monthly payment on a GHȼ10,000 loan at a 6% annual interest rate which is

amortized over 15years.

3. Find the monthly payments of an auto loan of GHȼ20,000 to be amortized over a 5-year

period at a rate of 9%.


1
Expert's answer
2021-03-31T13:00:33-0400

a)



where

  • A = payment Amount per month
  • P = initial Principal (loan amount)
  • r = interest rate per month
  • n = total number of months

Monthly payments = 78,000 *{[0.096/12(1+0.096/12)^30*12]/[(1+0.096/12)^30*12-1])

=78,000*0.00848

= 661.44


b)Total amount Paid = 661.44*30*12

= 238,118.40

Total interest paid = 238,118.40 - 78,000

=160,118.40


c) Using the loan amortization formula above,

Monthly payment = 10,000 * {[0.006/12(1+0.006/12)^15*12]/((1+ 0.006/12)^15*12 - 1)

= 58.1914


d) Using the loan amortization formula above,

Monthly payment = 20,000 *{[0.009/12(1+ 0.009/12)^5*12]/(1+0.009/12)^5*12-1)

= 341.087